OpenStudy (itsonlycdeee):

The equation below gives the height h of a passenger on a Ferris wheel at any time t during the ride to be h = 137 − 122 cos (π/10)t where h is given in feet and t is given in minutes. Use this equation to find the times at which a passenger will be 120 feet above the ground during the first revolution. I found the first time, which is about 4.5, but I'm confused on how to find the second time.

OpenStudy (anonymous):

are you familiar with the unit circle? the are 2 quadrants in the unit circle which cos(t) is positive

OpenStudy (itsonlycdeee):

Yes, the 1st quadrant and the 4th quadrant.

OpenStudy (anonymous):

is t inside or outside of the cos?

OpenStudy (itsonlycdeee):

Outside

OpenStudy (primeralph):

Should be inside.

OpenStudy (primeralph):

Well, most likely.

OpenStudy (anonymous):

are you sure?, because then h = 137 − 122 cos (π/10)t would be linear, meaning there is only 1 time h = 120

OpenStudy (itsonlycdeee):

Oh, then inside does make more sense.

OpenStudy (anonymous):

120 = 137 − 122 cos (π/10*t) -17 = -112 cos(π/10*t) 17/112 = cos(pi/10*t) \[\cos^{-1} (17/112) = \frac{ \pi }{ 10 }t\]

OpenStudy (itsonlycdeee):

Thank you I just figured it out!